# General Order Virtual Element Methods for Neumann Boundary Optimal Control Problems in Saddle Point Formulation

### v1

@@ 4. Error Estimates
On the other hand, the term ${\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varsigma}_{h}\right)-{\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)$ in (57) is estimated adding and subtracting ${\mathcal{A}}_{h}\left(\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)$, applying the definition of ${\mathcal{A}}_{h}$ and ${\mathcal{A}}$, the definition and the continuity of $\Pi^{0,E}_{k}$ together with the Cauchy-Schwarz inequality, a Poincarè inequality and (31), and (50), i.e.
(59) ${\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varsigma}_{h}\right)-{\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)$ $={\mathcal{A}}_{h}\left(\boldsymbol{x}_{h}-\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)+{\mathcal{A}}_{h}\left(\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)-{\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)$
$=\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(\Pi^{0,E}_{k}(y_{h}-y),\Pi^{0,E}_{k}w_{h}\right)_{E}$
@@
$\quad\quad+\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(\Pi^{0,E}_{k}y,\Pi^{0,E}_{k}w_{h}\right)_{E}-\left(y,w_{h}\right)_{E}$
@@
$\quad+\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y-\Pi^{0,E}_{k}y,\Pi^{0,E}_{k}w_{h}-w_{h}\right)_{E}$
$\lesssim\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{h}}\right\rVert_{X}\left\lVert{\boldsymbol{\varsigma}_{h}}\right\rVert_{X}+h^{s+1}\left\lvert{p}\right\rvert_{s+1}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}$
@@
5. Numerical Results
In this section, we present two numerical tests, that are solved with the proposed scheme using the dofi-dofi stabilization defined in Remark 2. The first test assess the numerical rate of convergence, while the second one explores the choice of the stability parameter $\sigma$. In the framework of the second test, we propose an alternative scheme designed to avoid issues related to the choice of $\sigma$.
5.1. Test 1
@@
$(0,0)$$(1,0)$$(1,1)$$(0,1)$$\Omega$$\Gamma_{D}$$\Gamma_{C}$
Figure 1. Test 1: schematic representation of the domain.
@@
be the exact solutions of Problem 3 with $\alpha=1$, $k(x)\equiv\gamma(x)\equiv 1$ and
$f(x_{1},x_{2})$ $=(1-x_{1})\sin(\pi x_{2})(1+\pi^{2})\,,$
$y_{\mathrm{d}}(x_{1},x_{2})$ $=\sin(\pi x_{2})\left[(1-x_{1})+\left(\frac{5\pi^{2}}{4}+1\right)\cos\left(\frac{\pi}{2}x_{1}\right)\right].$
@@
Figure 2. Test 1: meshes.
@@
Figure 3. Test 1: convergence plots for the state variable $y$.
@@
Figure 4. Test 1: convergence plots for the adjoint variable $p$.
@@
Figure 5. Test 1: convergence plots for the control variable $u$.
We solve Problem 4 with the stability parameter $\sigma=1$ for $k=1,\ldots,4$ on three different grids, depicted in Figure 2, obtaining the approximated solutions $y_{h},u_{h}$ and $p_{h}$. The considered meshes are: a tessellation of squares (named Cartesian, see Figure 2(a)), a mesh of non-convex octagons and nonagons (named Star, Figure 2(b)) and mesh of convex polygons obtained exploiting Polymesher [35] (named Polymesher, Figure 2(c)). For each mesh, we consider four refinement levels to show the convergence properties stated in Theorem 1, computing the following errors:
@@
    $\sqrt{\sum_{E\in\mathcal{M}_{h}}\left\lVert{\sqrt{\mathcal{K}}\left(\nabla\Pi^{\nabla,E}_{k}y_{h}-y\right)}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\left\lVert{\sqrt{\gamma}(\Pi^{0,E}_{k}y_{h}-y)}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}},$
@@
    $\sqrt{\sum_{E\in\mathcal{M}_{h}}\left\lVert{\sqrt{\mathcal{K}}\left(\nabla\Pi^{\nabla,E}_{k}p_{h}-p\right)}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\left\lVert{\sqrt{\gamma}(\Pi^{0,E}_{k}p_{h}-p)}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}}.$
@@
5.2. Test 2 (sensitivity study on the stabilization parameter $\sigma$)
$\Omega$$\Omega_{1}$$\Omega_{2}$$\Gamma_{D}$$\Gamma_{N}$$\Gamma_{C}$$\Gamma_{C}$$(0,0)$$(2,0)$$(2,1)$$(2,0)$
Figure 6. Test 2: schematic representation of the spatial domain.
(a) Polymesher60.
(b) Polymesher120.
Figure 7. Test 2: meshes.
(a) State.
(b) Adjoint.
(c) Control at $y=0$, $x\in[1,2]$.
(d) Control at $y=1$, $x\in[1,2]$.
Figure 8. Test 2: reference solution.
@@ 5.2. Test 2 (sensitivity study on the stabilization parameter $\sigma$)
Figure 9. Polymesher60. Errors with respect to reference solutions, for $\sigma=10^{-3},10^{-2},10^{1},1,10,100,1000$ and $k=1,2,3,4$.
(a) State, $L^{2}$ error.
(b) Adjoint, $L^{2}$ error.
(c) Control, $L^{2}$ error.
(d) State, e-error.
(e) Adjoint, e-error.
Figure 10. Polymesher120. Errors with respect to reference solutions, for $\sigma=10^{-3},10^{-2},10^{1},1,10,100,1000$ and $k=1,2,3,4$.
We propose a numerical study of VEM for several stabilization parameters $\sigma$, see Remark 2, for the following OCP on a setting inspired by [29]. The physical domain is $\Omega=(0,2)\times(0,1)$. The desired state is observed in $\Omega_{\text{obs}}=\Omega_{1}\cup\Omega_{2}$, for $\Omega_{1}=(1,2)\times(0.8,1)$ and $\Omega_{2}=(1,2)\times(0,0.2)$. The control boundary $\Gamma_{C}$ is $((1,2)\times\{1\})\cup((1,2)\times\{0\}).$ Homogeneous Neumann boundary conditions are imposed over $\{2\}\times(0,2)$. On $\Gamma_{D}=\partial\Omega\setminus(\Gamma_{C}\cup\Gamma_{N})$, we impose non-homogeneous Dirichlet boundary condition $g=1$. We remark that, even if the state equation features non-homogeneous Dirichlet boundary conditions, the adjoint equation still features homogeneous Dirichlet conditions, respectively [22]. The domain is represented in Figure 6. We solve Problem 4 with $\alpha=0.07$, $k(x)\equiv 1/12$, $\gamma(x)=1$, $f=0$ and $y_{\mathrm{d}}\equiv 2.5$ on two meshes, depicted in Figure 7: (i) Polymesher60, a conforming mesh with $60$ polygons built using Polymesher [35] and (ii) Polymesher120, another conforming Polymesher mesh of $120$ polygons. We solve the problem for $k=1,2,3,4$ and $\sigma=10^{-3},10^{-2},10^{-1},1,10,100,1000$. To test the accuracy of the VEM solution, we compare it to a reference finite element solution over a triangular mesh of $55000$ elements. The reference solution has been simulated though FEniCS library [26] and is depicted in Figure 8. From Figures 9 and 10, we can observe a variability of the errors both in term of L2 and energy norm. Let us focus on Figure 9, referring to the results obtained on Polymesher60. The minimum of the error in the L2 norm (top row) shifts from left to right increasing the polynomial degree. For example, if we consider state and adjoint and $k=1,2$ the optimal $\sigma$ is $10^{-1}$, while for $k=3,4$ the optimal value is $1$. For the control we see that $\sigma=1$ is optimal for $k=1,3,4$, while for $k=2$ the minimum L2 error is reached for $\sigma=10^{-1}$. From this, it is clear that different variables may give optimal error values for different $\sigma$. This is even more evident for the energy norm (bottom row) for state and adjoint for $k=2,3,4$: for the state the optimal value is $\sigma=1$, while for the adjoint variable smaller values of $\sigma$ would be beneficial. For $k=1$, the state optimal value in energy norm is reached for $\sigma=10^{-1}$, while for the adjoint is related to $\sigma=10^{3}$. Similar conclusions can be drawn from Figure 10, related to the results obtained on Polymesher120. We highlight that the optimal stabilization parameter may also vary in terms of the mesh we are using. For example, if we compare Figures 9(e) and 10(e) for $k=1$, the adjoint optimal error is reached for $\sigma=10^{3}$ and $\sigma=10^{-2}$, respectively.
5.3. Test 2 (structure preserving VEM)
In this section we tackle the numerical setting discussed in Section 5.2 through the employment of structure preserving VEM [9], to offer an alternative scheme that does not rely on an arbitrary non-polynomial stabilizing bilinear form. At the time of writing this paper, the scheme in [9] is available only for $k=1$. Under assumptions A.1 and A.2, for each $E\in\mathcal{M}_{h}$, let ${a}^{E}_{h}:\mathrm{H}^{1}\!(E)\times\mathrm{H}^{1}\!(E)\to\mathbb{R}$ be the local bilinear form defined as, for all $w_{h},q_{h}$ in $\mathrm{H}^{1}\!(E)$
(60) ${\tilde{a}}^{E}_{h}\left(w_{h},q_{h}\right)\vcentcolon=\left(\mathcal{K}\hat{\Pi}^{0,E}_{\ell_{E}}\nabla w_{h},\hat{\Pi}^{0,E}_{\ell_{E}}\nabla q_{h}\right)_{E}+\left(\gamma\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h},\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}q_{h}\right)_{E},$
where, given $\ell_{E}\in\mathbb{N}$, $\hat{\Pi}^{0,E}_{\ell_{E}}\nabla:\mathrm{H}^{1}\!(E)\rightarrow\boldsymbol{\mathrm{curl}}\,\mathbb{P}_{\ell_{E}+1}\!\left(E\right)$ is the $\mathrm{L}^{2}$-projection operator of functions in $\mathrm{H}^{1}\!(E)$, defined, for all $u\in\mathrm{H}^{1}\!(E)$, by the orthogonality condition
(61) $\left(\hat{\Pi}^{0,E}_{\ell_{E}}\nabla u,\boldsymbol{\mathrm{curl}}\,p\right)_{E}=\left(\nabla u,\boldsymbol{\mathrm{curl}}\,p\right)_{E}\quad\forall p\in\mathbb{P}_{\ell_{E}+1}\!\left(E\right)\,,$
with $\boldsymbol{\mathrm{curl}}\,p=\left(\frac{\partial p}{\partial y},-\frac{\partial p}{\partial x}\right)$ for any $p\in\mathbb{P}_{\ell_{E}+1}\!\left(E\right)$, and $\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}:\mathrm{H}^{1}\!(E)\to\mathbb{P}_{\ell_{E}+1}\!\left(E\right)$ is the operator defined, for all $u\in\mathrm{H}^{1}\!(E)$, by
(62) $\left(\boldsymbol{\mathrm{curl}}\,\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}u,\boldsymbol{\mathrm{curl}}\,p\right)_{E}$ $=\left(\nabla u,\boldsymbol{\mathrm{curl}}\,p\right)_{E}\quad\forall p\in\mathbb{P}_{\ell_{E}+1}\!\left(E\right)\,,$
$\int_{E}\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}u-u$ $=0\,.$
We remark that considering together the two projectors the following relation holds true
(63) $\left(\hat{\Pi}^{0,E}_{\ell_{E}}\nabla u-\boldsymbol{\mathrm{curl}}\,\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}u,\boldsymbol{\mathrm{curl}}\,p\right)_{E}=0\quad\forall p\in\mathbb{P}_{\ell_{E}+1}\!\left(E\right)\,,$
hence $\hat{\Pi}^{0,E}_{\ell_{E}}\nabla u=\boldsymbol{\mathrm{curl}}\,\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}u$ for any $u\in\mathrm{H}^{1}\!(E)$. On $\ell_{E}$ we assume:
  1. H.1
    $\ell_{E}$ is chosen such that there exists a set of degrees of freedom for $p\in\mathbb{P}_{\ell_{E}+1}\!\left(E\right)$ containing the integral mean $\frac{1}{\left\lvert{e}\right\rvert}\int_{e}p$ on each edge $e\in{\mathcal{E}_{E}}$.
We refer the reader to [9, Theorem 3] for conditions over $E$ to guarantee H.1. We here propose an extension of [9, Theorem 2].
Theorem 2.
Assuming A.1, A.2 and H.1, for every $E\in\mathcal{M}_{h}$, the discrete bilinear form (60) is coercive on $Y_{h}(E)$, i.e. there exists $\eta^{E}_{*}>0$, independent of $h_{E}$, such that
(64) ${\tilde{a}}^{E}_{h}\left(w_{h},w_{h}\right)\geq\eta^{E}_{*}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y(E)}^{2}\quad\forall w_{h}\in Y_{h}(E)\,.$
Proof.
Let $w_{h}\in Y_{h}(E)$ be given. Theorem 2 in [9] states that, under the current assumptions,
(65) $\left(\hat{\Pi}^{0,E}_{\ell_{E}}\nabla w_{h},\hat{\Pi}^{0,E}_{\ell_{E}}\nabla w_{h}\right)_{E}\gtrsim\left\lVert{\nabla w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}\,.$
Now, we want to show that $\left\lVert{\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}u}\right\rVert_{\mathrm{L}^{2}\!(E)}\gtrsim\left\lVert{u}\right\rVert_{\mathrm{L}^{2}\!(E)}$. Considering the definition of the projectors $\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}$ and $\hat{\Pi}^{0,E}_{\ell_{E}}\nabla$ and (63), $\left\lVert{\nabla\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}=\left\lVert{\hat{\Pi}^{0,E}_{\ell_{E}}\nabla w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}$ holds true. Hence, we can infer that $\left\lVert{\nabla\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}\gtrsim\left\lVert{\nabla w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}$.
Let us recall that
$\Pi^{0,E}_{0}w_{h}=\left\lvert{E}\right\rvert^{-1}\int_{E}w_{h}=\left\lvert{E}\right\rvert^{-1}\int_{E}\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}\,,$
where the last equality is due to (62). It implies that
$\left\lVert{\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}$ $=\left\lVert{\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}-\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\left\lVert{\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2},$
$\left\lVert{w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}$ $=\left\lVert{w_{h}-\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\left\lVert{\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}\,.$
Moreover, notice that the following Poincaré inequality
$\left\lVert{w_{h}-\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}\lesssim h_{E}\left\lVert{\nabla w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}$
and the inverse inequality
$\left\lVert{\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}-\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}\gtrsim h_{E}\left\lVert{\nabla\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}$
hold true. Hence, we obtain
(66) $\left\lVert{\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}$ $=\left\lVert{\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}-\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\left\lVert{\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}$
$\gtrsim h^{2}_{E}\left\lVert{\nabla\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\left\lVert{\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}$
$\gtrsim h_{E}^{2}\left\lVert{\nabla w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\left\lVert{\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}$
$\gtrsim\left\lVert{w_{h}-\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\left\lVert{\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}\gtrsim\left\lVert{w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}\,.$
Hence, $\left\lVert{\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}$ is a norm for each $w_{h}\in Y_{h}(E)$. Finally, under the hypothesis on $\mathcal{K}$ and $\gamma$, it is easy to see that (65) and (66) imply that the thesis holds true. ∎
@@ 5.3. Test 2 (structure preserving VEM)
Figure 11. Polymesher60. Errors with respect to reference solutions, for $\sigma=10^{-3},10^{-2},10^{1},1,10,100,1000$ and $k=1,2,3,4$.
(a) State, $L^{2}$ error.
(b) Adjoint, $L^{2}$ error.
(c) Control, $L^{2}$ error.
(d) State, e-error.
(e) Adjoint, e-error.
Figure 12. Polymesher120. Errors with respect to reference solutions, for $\sigma=10^{-3},10^{-2},10^{1},1,10,100,1000$ and $k=1,2,3,4$.
By Theorem 2, we can solve Problem 4 substituting for each $E\in\mathcal{M}_{h}$ ${a}^{E}_{h}$ with ${\tilde{a}}^{E}_{h}$ and retrieve the well-posedness of the method (as in Proposition 1) and the error estimates (as in Theorem 1). In practice, for the computation of ${\tilde{a}}^{E}_{h}$ we exploit the QR-based algorithm presented in [9], that returns the matrix $\boldsymbol{\Pi}=\left\{\pi_{i,j}\right\}_{ij}$ such that
(67) $\hat{\Pi}^{0,E}_{\ell_{E}}\nabla\phi_{j}=\sum_{i=1}^{\dim\mathbb{P}_{\ell_{E}+1}\!\left(E\right)-1}\pi_{ij}\,\boldsymbol{\mathrm{curl}}\,m_{i}\,,\quad\forall j=1,\ldots,N_{E}$
where $\{\phi_{j}\}_{j=1}^{N_{E}}$ is a set of basis functions of $Y_{h}(E)$ and $\{m_{i}\}_{i=1}^{\dim\mathbb{P}_{\ell_{E}+1}\!\left(E\right)}$ is a set of basis functions of $\mathbb{P}^{0}_{\ell_{E}+1}\!\left(E\right)\vcentcolon=\left\{p\in\mathbb{P}_{\ell_{E}+1}\!\left(E\right):\int_{E}p=0\right\}$. Then, using (63) we get
(68) $\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}\phi_{j}=\left\lvert{E}\right\rvert^{-1}\int_{E}\phi_{j}+\sum_{i}\pi_{ij}\,m_{i}\,.$
Hence, the computation of the mass term in ${\tilde{a}}^{E}_{h}$ does not require an additional computational cost.
Now, we present the results obtained with this method. We compare the errors with the ones discussed in Section 5.2 for $k=1$. From Figures 11 and 12 we see that the errors made by structure-preserving VEM (named stab-free in legends) are quite close to the optimal errors made by stabilized VEM, independently of the solution and of the mesh. This suggests that this method is a valid alternative to avoid issues related to the choice of the stability parameter.
6. Conclusions
In this work, we focus on the VEM approximation of Neumann boundary OCPs. We address the problem in saddle point formulation, providing a rigorous a priori error estimation for all the variables at hand. To the best of our knowledge, this is the first time that the analysis is proposed in the saddle point framework for arbitrary polynomial degrees. A numerical test based on analytic solutions confirmed the expected theoretical convergence rates. A second test is proposed: we show the accuracy of the method over a ground truth solution based on FEM discretization. For this latter experiment, we analyze the role of the VEM stabilization and its influence on the approximation error of each component for this coupled problem. We observe mesh-dependency effects and variabilities between the three variables of the system. To circumvent this issue we also propose a structure preserving strategy for polynomial degree $k=1$.
This contribution is meant to be a bridge between the theoretical and applied aspects of the problem, where no exact solution is known. Indeed, this work is a first step towards more complex applications. Indeed, there are many aspects to be investigated to better show the VEM potential for such coupled problems, for example more complicated geometrical structures and models. Moreover, a further interesting potential development would be the extension of theoretical and numerical analysis to other type of OCPs, with different control boundary conditions and constraints.

### v2

@@ 4 Error Estimates
On the other hand, the term ${\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varsigma}_{h}\right)-{\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)$ in (67) is estimated adding and subtracting ${\mathcal{A}}_{h}\left(\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)$, applying the definition of ${\mathcal{A}}_{h}$ and $\mathcal{A}$}, the definition and the continuity of $\Pi^{0,E}_{k}$ together with the Cauchy-Schwarz inequality, a Poincaré inequality and (39), and (61), i.e.
${\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varsigma}_{h}\right)-{\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)$ $={\mathcal{A}}_{h}\left(\boldsymbol{x}_{h}-\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)+{\mathcal{A}}_{h}\left(\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)-{\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)$ (69)
$=\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left(\Pi^{0,E}_{k}(y_{h}-y),\Pi^{0,E}_{k}w_{h}\right)_{E}$
@@
$\quad\quad+\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left(\Pi^{0,E}_{k}y,\Pi^{0,E}_{k}w_{h}\right)_{E}-\left(y,w_{h}\right)_{E}$
@@
$\quad+\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left(y-\Pi^{0,E}_{k}y,\Pi^{0,E}_{k}w_{h}-w_{h}\right)_{E}$
$\lesssim\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{h}}\right\rVert_{X}\left\lVert{\boldsymbol{\varsigma}_{h}}\right\rVert_{X}+h^{s+1}\left\lvert{y}\right\rvert_{s+1}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}$
@@
Remark 3 (Standard VEM approximation).
Substituting $a_{h}$ with $\tilde{a}_{h}$ defined in (28), Theorem 1 still holds true, following [3] for the VEM counterpart of estimates (48) and (65). To the best of the authors’ knowledge, the complete a priori error estimates of a Neumann boundary optimal control problem in a saddle point framework, have never been addressed before in the VEM community.
5 Numerical Results
In this section, we present three numerical experiments designed to assess the stability and robustness of the proposed method. The first test is a convergence study aimed at validating the theoretical results established in the previous section. The second test examines the dependence of the standard VEM procedure—described in Remark 2—on the choice of the stabilization parameter $\sigma$, and compares its accuracy with that of the proposed method. Finally, the third test introduces a more application‑oriented setting to assess the robustness of the proposed approach.
5.1 Test 1
@@
[…]
Figure 1: Test 1: schematic representation of the domain.
@@
be the exact solutions of Problem 3 with $\alpha=1$, $\mathcal{K}(x)\equiv\gamma(x)\equiv 1$, $\beta=[0,1]$ and
$f(x_{1},x_{2})$ $=(1-x_{1})\left[\sin(\pi x_{2})(1+\pi^{2})+\pi\cos(\pi x_{2})\right]\,$,
$y_{\mathrm{d}}(x_{1},x_{2})$ $=\sin(\pi x_{2})\left[(1-x_{1})+\left(\frac{5\pi^{2}}{4}+1\right)\cos\left(\frac{\pi}{2}x_{1}\right)\right]-\pi\cos(\pi x_{2})\cos\left(\frac{\pi}{2}x_{1}\right).$
@@
Figure 2: Test 1: meshes.
@@
Figure 3: Test 1: convergence plots for the state variable $y$.
@@
Figure 4: Test 1: convergence plots for the adjoint variable pp.
@@
Figure 5: Test 1: convergence plots for the control variable $u$.
We solve Problem 4 for $k=1,\ldots,4$ on three different grids, depicted in Figure 2, obtaining the approximated solutions $y_{h},u_{h}$ and $p_{h}$. The considered meshes are: a tessellation of squares (named Cartesian, see Figure 2(a)), a mesh of non-convex octagons and nonagons (named Star, Figure 2(b)) and mesh of convex polygons obtained exploiting Polymesher [31] (named Polymesher, Figure 2(c)). For each mesh, we consider four refinement levels to show the convergence properties stated in Theorem 1, computing the following errors:
@@
    $\sqrt{\sum_{E\in\mathcal{M}_{h}}\left\lVert{\sqrt{\mathcal{K}}\left(\nabla\Pi^{\nabla,E}_{k}y_{h}-y\right)}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}},$
@@
    $\sqrt{\sum_{E\in\mathcal{M}_{h}}\left\lVert{\sqrt{\mathcal{K}}\left(\nabla\Pi^{\nabla,E}_{k}p_{h}-p\right)}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}}.$
@@
5.2 Test 2 (sensitivity study on the stabilization parameter $\sigma$)
@@
Figure 6: Star. Errors with respect to reference solutions, for $\sigma=10^{-3},10^{-2},10^{-1},1,10,100,1000$ and $k=1,2$.
@@
Figure 7: Polymesher. Errors with respect to reference solutions, for $\sigma=10^{-3},10^{-2},10^{-1},1,10,100,1000$ and $k=1,2$.
In this section we consider the same problem as in Section 5.1 and perform a numerical study of the errors done by the stabilized VEM scheme described in Remark 2 for several choices of stabilization parameter $\sigma\in\{10^{-3},10^{-2},10^{1},1,10,100,1000\}$. These results are compared with the results of the proposed stabilization-free scheme. In particular, we consider the most refined meshes belonging to the families Star and Polymesher depicted in Figure 2, both featuring around 3000 polygons. In Figures 6 and 7 we display the $L^{2}$ and energy errors corresponding to the discrete solutions for $k=1,2$. We can observe a significant dependence of the error of stabilized VEM on the stabilization parameter. Moreover, the values of $\sigma$ corresponding to minima depend on the mesh (see in particular the results for the control variable). Finally, we observe that the errors made by the proposed stabilization-free scheme are always very close to the optimal errors made by stabilized VEM, independently of the solution and of the mesh. This suggests that the proposed method is a valid alternative to avoid issues related to the tuning of the stabilization parameter.
5.3 Test 3 (towards applications)
$\Omega$$\Omega_{1}$$\Omega_{2}$$\Gamma_{D}$$\Gamma_{N}$$\Gamma_{C}$$\Gamma_{C}$(0,0)(0,0)(2,0)(2,0)(2,1)(2,1)$(0,2)$
Figure 8: Test 3: schematic representation of the spatial domain.
Figure 9: Test 3: Comparison of state and adjoint solutions obtained with the proposed method and the linear finite element implemented in FEniCS.
(a) Control at $y=0$, $x\in[1,2]$.
(b) Control at $y=1$, $x\in[1,2]$.
Figure 10: Test 3: reference solution.
We propose a numerical study of the robustness of the proposed method, in its lowest-order formulation, by comparing it with a linear Finite Element discretization performed by the FEniCS library [21] on the following OCP, in a setting inspired by [25]. The physical domain is $\Omega=(0,2)\times(0,1)$. The desired state is observed in $\Omega_{\text{obs}}=\Omega_{1}\cup\Omega_{2}$, for $\Omega_{1}=(1,2)\times(0.8,1)$ and $\Omega_{2}=(1,2)\times(0,0.2)$. The control boundary $\Gamma_{C}$ is $((1,2)\times\{1\})\cup((1,2)\times\{0\}).$ Homogeneous Neumann boundary conditions are imposed over $\{2\}\times(0,2)$. On $\Gamma_{D}=\partial\Omega\setminus(\Gamma_{C}\cup\Gamma_{N})$, we impose non-homogeneous Dirichlet boundary condition $g=1$. We remark that, even if the state equation features non-homogeneous Dirichlet boundary conditions, the adjoint equation still features homogeneous Dirichlet conditions, see [22]. The setting of this problem is represented in Figure 8. We solve Problem 4 with $\alpha=0.07$, $\mathcal{K}=1/12$, $\gamma=1$, $\beta=[1,1]$, $f=0$ and $y_{\mathrm{d}}\equiv 2.5$ on a conforming (w.r.t $\Omega_{\text{obs}}$) mesh built using Polymesher [31], with $5490$ points. We compare the SFVEM solution to the FEniCS reference FEM solution over a triangular mesh with $28130$ points. In Figure 9, we report the results of the state and the adjoint variables obtained by the SFVEM method and the FEM method. The SFVEM solution matches the FEM reference. Figure 10 shows the comparison of the solutions related to the control. The SFVEM control follows the behavior of the FEM reference, except for some small numerical oscillations around $x=1$ due to the steep gradients in the solutions. This limitation can be effectively mitigated by refining the computational mesh in the region of interest. A more efficient strategy would involve an adaptive mesh refinement procedure, guided by an appropriate a posteriori error estimator. The development of this adaptive framework represent a promising direction for a future work.
6 Conclusions
In this work, we focused on a Stabilization-Free VEM approximation of Neumann boundary OCPs. We addressed the problem in saddle point formulation, providing a rigorous a priori error estimation for all the variables at hand. A numerical test based on analytic solutions confirmed the expected theoretical convergence rates. In a second test, we analyzed the role of the stabilization term in the standard VEM formulation and its influence on the approximation error of each component for this coupled problem, comparing it with the error made by the SFVEM scheme. For standard VEM, we observed mesh-dependency effects and variabilities between the three variables of the system, which are well circumvented by the SFVEM approach. In the last test, we showed the accuracy and robustness of the proposed method on a more application-oriented test. This last test is meant to be a bridge between the theoretical and applied aspects of the problem, where no exact solution is known. Indeed, this work is a first step towards more complex applications. There are many aspects to be investigated to better show the SFVEM potential for such coupled problems, for instance, more complicated geometrical structures and models, and the development of adaptive strategies.